ASAP. Please show all work thank you

The average (mean) value of ?(?) is given by

1/ ?−? ∫( b to a)?(?)??

on the interval (a,b) and is denoted by ????rage. Dont use another interval

(a) Find the average value of any quadratic function with vertex at the origin on the interval [?, ?].

(b) Find a number ? in the given interval such that ?(?) = ???erage
Answer #1

Key point here is, any quadratic with vertex at origin is of the form f(x)=kx^2

proof. Let f(x)= ax^2+bx+c, since f(0)=0 (it passes through origin) gives c=0. also slope at (0,0) is 0. f'(x)= 2ax+b putting f'(x) zero at zero gives b=0. hence f is of the form ax^2.

now the proof is easy

b qurage (mean value of f(x) is given by f f(x) dx on the internal [a, b] - ba a quadratic function with verlex at [0;o) look